OBSERVATION OF C−H⋯π INTERACTIONS: MICROWAVE SPECTRA AND STRUCTURES OF THE CH_{2}FX⋯HCCH (X=F,Cl) WEAKLY BOUND COMPLEXES

LENA F. ELMUTI, \underline{DANIEL A. OBENCHAIN}, DON L. JURKOWSKI, AMELIA J. SANDERS, REBECCA A. PEEBLES, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Avenue, Charleston, IL 61920; AMANDA L. STEBER, JUSTIN L. NEILL, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., PO Box 400319, Charlottesville, VA 22904.

With an interest in characterizing C−H⋯π interactions, CH_{2}F_{2}⋯HCCH and CH_{2}ClF⋯HCCH have been examined by Fourier-transform microwave (FTMW) spectroscopy. These interactions involve the π bond in acetylene acting as a hydrogen bond acceptor to both hydrogen atoms of CH_{2}FX. In addition, there is a secondary contact between one hydrogen atom from acetylene and the X atom in the halomethane (X=F in CH_{2}F_{2}, X=Cl in CH_{2}ClF).

Initial assignments for the most abundant isotopologues of both species were completed using the chirped-pulse FTMW spectrometers at the University of Virginia (CH_{2}ClF⋯HCCH) and at Eastern Illinois University (CH_{2}F_{2}⋯HCCH). Rotational constants obtained from experiment are in good agreement with those of the most stable orientations predicted by ab initio calculations at the MP2/6-311++G(2d,2p) level. Multiple isotopically substituted species for each complex were measured using a Balle-Flygare cavity FTMW spectrometer at Eastern Illinois University. Spectroscopic parameters for all observed isotopologues will be presented, and a comparison of the C−H⋯ π interactions in these and related complexes will be discussed.